HUBBLE SPACE TELESCOPE DAILY REPORT #2580 PERIOD COVERED: 0000Z (UTC) 03/14/00 - 0000Z (UTC) 03/15/00 Daily Status Report as of 075/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 8460 (Cycle 8 Supplemental Darks pt2/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations completed with no reported problems. 1.2 Completed Three Sets of STIS/CCD/MA1 8038 (Imaging of the Alpha Orionis System) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make ultraviolet images of the Alpha Orionis triple system using different FUV/NUV filters. The goal is to resolve surface features and any extended atmosphere. Images will also be recorded of a reference point source star in order to determine exactly the PSF for the Alpha Orionis observations. The observations completed with no reported problems. 1.3 Completed STIS/CCD/MA2 8421 (Spectroscopic Sensitivity for (MAMA and) CCD) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to perform basic sensitivity measurement for some supported (MAMA and) CCD spectroscopic modes. Sensitivity measurements are done for a few primary tilts of the gratings. The proposal completed nominally, with no reported problems. 1.4 Completed Three Sets of WF/PC-2 8442 (WF/PC-2 Cycle 8 Standard Darks) The WF/PC-2 was used to make a series of Cycle 8 standard dark calibration observations. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed STIS/CCD 8439 (Bias Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no reported problems. 1.6 Completed Two Sets of STIS/CCD 8437 (Dark Monitor-Part 2 C8) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. The proposal completed nominally. 1.7 Completed WF/PC-2 8239 (Imaging the NLR in a Complete Sample of z<0.5 Radio-Quiet PG Quasars) The WF/PC-2 was used to observe quasar PG1307+085 and to determine its role in the evolution of galaxies, where a central engine leaves its visible imprints in the galaxy through strong ionization and heating of the ISM. The observations completed with no reported problems. 1.8 Completed STIS/CCD 8786 (Morphological and Spectral Evolution of Galaxies: Follow-up on NGC346-1) The Space Telescope Imaging Spectrograph (CCD) was used to observe the high-redhsift universe by providing kiloparsec resolution of distant galaxies. However, since the rest-frame wavelengths of galaxies move further into the UV at increasingly higher redshifts, evolutionary effects are diluted by insufficiently well determined k-corrections. Thus a library of UV images of normal galaxies at low redshifts is essential to our understanding of the properties and evolution of the high redshift universe. The proposal completed nominally, with no reported problems. 1.9 Completed STIS/CCD 8253 (Jet Acceleration of Narrow Line Region Gas) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the Markarian galaxy MARK463E to study its jet acceleration of narrow line Gas regions. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed STIS/CCD 8412 (STIS CCD Imaging Flats C8) The Space Telescope Imaging Spectrograph (CCD) was used to investigate the flat-field stability of the unit over a monthly period. A bias calibration observation was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed FGS 8525 (FGS-2R Guider Commissioning) FGS-2R was used to verify the acquisition logic for the newly-replace FGS, including K factors, alignment, photometric calibration and the ability to lock on both bright and faint guide stars. There were no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 17 Successful: 17 Scheduled Re-acquisitions: 5 Successful: 5 2.2 FHST Updates: Scheduled: 21 Successful: 21 2.3 Operations Notes: A TTR was written when a GCMR could not be sent at 074/1825Z due to a problem with the NCC software. Per ROP DF-18A, the ESB was dumped and cleared at 074/1914Z. There was a STIS EMC retry at 075/0644Z. The STIS flight software error counter was reset. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW